OpenStudy (anonymous):

8x - 1 = -4x + 23 x =

OpenStudy (anonymous):

8x - 1 = -4x + 23 12x - 1 = 23 12x = 24 x=2

OpenStudy (anonymous):

8x-1=-4x+23 add +4x to both sides 8x-1+4x=23 add 8x+4x=12x 12x-1=23 add +1 to both sides 12x=24 Divide both sides by 12 x=2

OpenStudy (anonymous):

6x - 8 + 2x - 5 = 7 - 2x x =

OpenStudy (anonymous):

(open a new question)

OpenStudy (anonymous):

6x-8+2x-5=7-2x add 2x to both sides 6x-8+2x+2x-5=7 add -8+-5=-13 add 6x+2x+2x=10x 10x-13=7 add 13 to both sides 10x=20 Divide by 10 on both sides x=2